# HWA-UNETR: Hierarchical Window Aggregate UNETR for 3D Multimodal Gastric Lesion Segmentation

### v2

@@ 1 Introduction
Gastric Cancer (GC) ranks as the fifth most common malignant neoplasm globally and the fourth leading cause of cancer-related mortality [20], whose detection of lesions is a critical step in clinical treatment. Through sustained technical refinements, Magnetic Resonance Imaging (MRI) has evolved into a superior modality for gastrointestinal malignancy evaluation [4], with its diagnostic ascendancy over Computed Tomography (CT) stemming from three cardinal virtues: multi-parametric soft tissue discrimination, absence of ionizing radiation, and synergistic incorporation of functional molecular imaging paradigms [19]. Recently, with the advancement of Deep Learning (DL) technologies, quantitative multimodal MRI scans have leveraged their inherent advantages to serve as an essential tool for detecting cancerous lesions [22, 8]. More importantly, these technologies have significantly enhanced medical image segmentation methods [16, 11], allowing for the systematic delineation of malignant tumor spatial distributions and precise quantification of lesion heterogeneity through the integration of multimodal data.
Existing DL methods are sensitive to imaging modalities and lesion characteristics, leading to segmentation performance that relies heavily on training data quality [3]. However, the scarcity of large-scale, high-quality multimodal MRI datasets for GC necessitate reliance on generic medical imaging datasets and transfer learning [14], both of which are resource-intensive and reduce specificity. Furthermore, gastric cancer lesion detection involves multiple spatially misaligned MRI modalities, which pose two critical challenges for DL methods: (1) aligning inherently misaligned imaging modalities and integrating complementary features, and (2) fully leveraging the complementary characteristics of each modality for multi-scale context modeling. Previous studies enhance segmentation via feature fusion [21, 23], yet clinical applicability remains limited due to reliance on overly optimistic data assumptions. Additionally, improving model representational capacity has proven effective for segmentation tasks. CNN-based [10, 17] and Transformer-based methods [8, 11, 16] improve contextual feature modeling, while Mamba-based methods [22, 6] optimize the modeling process by incorporating state space models (SSMs). Nonetheless, these predominantly focus on pixel-level predictions within single modalities, neglecting multimodal misalignment and complementarity. Recently, H-DenseFormer [18] proposes a multi-path parallel embedding framework to enhance segmentation efficiency and accuracy in multimodal tasks. Meanwhile, MMEF [5] integrates Dempster-Shafer theory [5] with a DL network to achieve reliable feature extraction. Nevertheless, existing approaches still struggle to fully leverage the complementary features of gastric MRI modalities, limiting their effectiveness in achieving reliable lesion segmentation.
@@ 2 GCM 2025: Gastric Cancer MRI Dataset
Dataset Construction: MRI scans of GCM 2025 were acquired from November 2017 to December 2024. All examinations were performed using 1.5T or 3.0T scanners equipped with dedicated body coils. Sensitive patient information in this dataset has been removed. Each volume was annotated by a professional doctor and cross-checked by a second doctor. Detailed data acquisition information are provided at this public https URL.
@@ 3.1 U-shaped Framework Overview
HWA-UNETR adopts a 4-stage U-shaped architecture comprising an encoder-decoder structure with skip connections, which is shown in Fig. 2 (a) and (b), aiming to achieve high-performance 3D multimodal image segmentation. The input volumetric data $X_{in}$ with dimensions $H_{in}\times W_{in}\times D_{in}\times C_{in}$ (height, width, depth, modals) undergoes hierarchical feature extraction in the encoder. At each stage $i$, spatial resolution is progressively reduced to $\frac{H}{2^{i+1}}\times\frac{W}{2^{i+1}}\times\frac{D}{2^{i+1}}$ via depthwise separable convolutional layers followed by instance normalization and ReLU activation. In the decoder, transposed convolutional layers upsample feature maps to recover the original spatial resolution. Simultaneously, skip connections integrate multi-scale encoder features to minimize information loss. Finally, a segmentation head equipped with a sigmoid activation layer produces pixel-wise probability maps, facilitating precise and scalable image segmentation.
@@ 3.2 Hierarchical Window Aggregate Block
The HWA block, whose structure is shown in Fig. 2 (c), operates at the multimodal volumetric input $X_{in}$ stage, dynamic window partitioning, and multi-kernel convolutions are employed to align cross-modal local semantics. This structure allows for the extraction of features from multimodal images with cross-resolution characteristics, generating semantically consistent fused image features. Specifically, the HWA Block adaptively employs depthwise convolutional layers with varying strides $r$ and kernel sizes $k$ based on the input image shapes to extract features from multimodal images, as detailed in Eq. (1).
@@ 3.3 Stratified Group Convolution Block
The SGC block, deployed post-downsampling in each encoder stage, enhances local feature retention and spatial information preservation for cross-level fusion [22]. As shown in Fig. 2 (d), the tri-branch architecture processes shared input $X_{si}$ via: (1) dual depthwise convolutions ($k$=3,1) with instance normalization obtain $X_{1}$; (2) pointwise convolution refinement obtain $X_{2}$; and (3) residual-enhanced fusion with MLP transformation obtain $X_{so}$ as result of SGC block.
@@ 3.4 Tri-orientated Fusion Mamba Block
The TFM Block is deployed in the skip connections stage, specifically designed for modeling global and multi-scale features, as shown in Fig. 2 (e).

### v3

@@ 1 Introduction
Gastric Cancer (GC) ranks as the fifth most common malignant neoplasm globally and the fourth leading cause of cancer-related mortality [23], whose detection of lesions is a critical step in clinical treatment. Through sustained technical refinements, Magnetic Resonance Imaging (MRI) has evolved into a superior modality for gastrointestinal malignancy evaluation [5], with its diagnostic ascendancy over Computed Tomography (CT) stemming from three cardinal virtues: multi-parametric soft tissue discrimination, absence of ionizing radiation, and synergistic incorporation of functional molecular imaging paradigms [22]. Recently, with the advancement of Deep Learning (DL) technologies, quantitative multimodal MRI scans have leveraged their inherent advantages to serve as an essential tool for detecting cancerous lesions [25, 12, 9]. More importantly, these technologies have significantly enhanced medical image segmentation methods [19, 13], allowing for the systematic delineation of malignant tumor spatial distributions and precise quantification of lesion heterogeneity through the integration of multimodal data.
Existing DL methods are sensitive to imaging modalities and lesion characteristics, leading to segmentation performance that relies heavily on training data quality [18, 4, 1]. However, the scarcity of large-scale, high-quality multimodal MRI datasets for GC necessitate reliance on generic medical imaging datasets and transfer learning [16], both of which are resource-intensive and reduce specificity. Furthermore, gastric cancer lesion detection involves multiple spatially misaligned MRI modalities, which pose two critical challenges for DL methods: (1) aligning inherently misaligned imaging modalities and integrating complementary features, and (2) fully leveraging the complementary characteristics of each modality for multi-scale context modeling. Previous studies enhance segmentation via feature fusion [24, 26], yet clinical applicability remains limited due to reliance on overly optimistic data assumptions. Additionally, improving model representational capacity has proven effective for segmentation tasks. CNN-based [11, 20] and Transformer-based methods [9, 13, 19] improve contextual feature modeling, while Mamba-based methods [25, 7] optimize the modeling process by incorporating state space models (SSMs). Nonetheless, these predominantly focus on pixel-level predictions within single modalities, neglecting multimodal misalignment and complementarity. Recently, H-DenseFormer [21] proposes a multi-path parallel embedding framework to enhance segmentation efficiency and accuracy in multimodal tasks. Meanwhile, MMEF [6] integrates Dempster-Shafer theory [6] with a DL network to achieve reliable feature extraction. Nevertheless, existing approaches still struggle to fully leverage the complementary features of gastric MRI modalities, limiting their effectiveness in achieving reliable lesion segmentation.
@@ 2 GCM 2025: Gastric Cancer MRI Dataset
Dataset Construction: MRI scans of GCM 2025 were acquired from November 2017 to December 2024. All examinations were performed using 1.5T or 3.0T scanners equipped with dedicated body coils. Sensitive patient information in this dataset has been removed. Each volume was annotated by a professional doctor and cross-checked by a second doctor. Detailed data acquisition information is provided at this public https URL.
@@ 3.1 U-shaped Framework Overview
HWA-UNETR adopts a 4-stage U-shaped architecture comprising an encoder-decoder structure with skip connections, which is shown in Fig. 2 (a), aiming to achieve high-performance 3D multimodal image segmentation. The input volumetric data $X_{in}$ with dimensions $H_{in}\times W_{in}\times D_{in}\times C_{in}$ (height, width, depth, modals) undergoes hierarchical feature extraction in the encoder. At each stage $i$, spatial resolution is progressively reduced to $\frac{H}{2^{i+1}}\times\frac{W}{2^{i+1}}\times\frac{D}{2^{i+1}}$ via depthwise separable convolutional layers followed by instance normalization and ReLU activation. In the decoder, transposed convolutional layers upsample feature maps to recover the original spatial resolution. Simultaneously, skip connections integrate multi-scale encoder features to minimize information loss. Finally, a segmentation head equipped with a sigmoid activation layer produces pixel-wise probability maps, facilitating precise and scalable image segmentation.
@@ 3.2 Hierarchical Window Aggregate Block
The HWA block, whose structure is shown in Fig. 2 (b), operates at the multimodal volumetric input $X_{in}$ stage, dynamic window partitioning, and multi-kernel convolutions are employed to align cross-modal local semantics. This structure allows for the extraction of features from multimodal images with cross-resolution characteristics, generating semantically consistent fused image features. Specifically, the HWA Block adaptively employs depthwise convolutional layers with varying strides $r$ and kernel sizes $k$ based on the input image shapes to extract features from multimodal images, as detailed in Eq. (1).
@@ 3.3 Stratified Group Convolution Block
The SGC block, deployed post-downsampling in each encoder stage, enhances local feature retention and spatial information preservation for cross-level fusion [25]. As shown in Fig. 2 (c), the tri-branch architecture processes shared input $X_{si}$ via: (1) dual depthwise convolutions ($k$=3,1) with instance normalization obtain $X_{1}$; (2) pointwise convolution refinement obtain $X_{2}$; and (3) residual-enhanced fusion with MLP transformation obtain $X_{so}$ as result of SGC block.
@@ 3.4 Tri-orientated Fusion Mamba Block
The TFM Block is deployed in the skip connections stage, specifically designed for modeling global and multi-scale features, as shown in Fig. 2 (d).